08 July 2014 UT Mini-queue night

Summary

Night delayed due to heavy, threatening clouds and rain in the early evening. Finally opened at around 07:45 UT with some cloud cover at horizon but obvious holes at zenith. Had some alarms on the chiller that slowed us down; tried LBC on an OSU assassin but found that temperature difference was too large to collimate successfully (in small part because we had to turn off mirror ventilation during chiller issue) . Switched to MODS and started with v2468cyg which was successful (!) in some fairly good conditions (light cirrus and ~1" seeing) and ASASSN14dc target (light cirrus ~1" seeing). Overall a few good hours of observing.

Plan

Start with LBC (OSURC monitor program; failed); continued with MODS (v2468cyg; SUCCESS!); MODS (ASASSN14dc; SUCCESS!); MODS (specphot: BD284211)

Log

07:50 - Clouds are moving off; trying first to collimate at N6946d

07:52 - Running dofpia but need some initial coma

07:55 - Trying to add some coma with AO-by-eye

07:57 - Running dofpia again, much better fit for coma, but...

08:00 - Still struggling -- too much spherical; temperature is just too far off( because of closed dome and the chiller issues)

08:05 - Switching to MODS

08:25 - Finished switch and collimate

08:30 - Realized that fake guide star had been left in beam; calling JHill for assistance. Now we know how to reset this ourselves.

08:45 - Slewing to V2468Cyg

08:48 - Acquiring V2468Cyg

09:15 - Observing V2468Cyg; FWHM = 1.0. Some cirrus but sky has improved dramatically!

09:28 - Obvious emission lines so we have the right target; 0.82" seeing. Starting second spectra

09:39 - Observing one more spectra for V2468Cyg

09:53 - Slewing to ASASSN14dc; pointing is off. Adjust pointing and collimation

10:05 - Slewing to ASASSN14dc again

10:20 - Observing ASASSN14dc; FWHM = 1.0 from GCS

First image done before twilight; second image done while twilight occurs (10:39 about 200s into exposure); third exposure done during

twilight but still quite bright continuum compared to BG

10:57 - Slewing to specphot standard BD284211

11:10 - Observing specphot standard BD284211

11:27 - Closing; Running morning cals
Topic revision: r1 - 08 Jul 2014, MichelleEdwards
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback